A wheel with mass m = 1 kg under the action of a moment of 3 Nm, rotates at an angular acceleration of 0.5 rad / s². Find the radius of the wheel.


1
Expert's answer
2021-04-01T18:42:34-0400

We can find the radius of the wheel as follows:


"\tau=I\alpha,""\tau=\dfrac{1}{2}mr^2\alpha,""r=\sqrt{\dfrac{2\tau}{m\alpha}}=\sqrt{\dfrac{2\cdot3\ N\cdot m}{1\ kg\cdot0.5\ \dfrac{rad}{s^2}}}=3.5\ m."